# Reinforcement Learning from Bagged Reward

### v2

@@ Abstract
In Reinforcement Learning (RL), it is commonly assumed that an immediate reward signal is generated for each action taken by the agent, helping the agent maximize cumulative rewards to obtain the optimal policy. However, in many real-world scenarios, immediate reward signals are not obtainable; instead, agents receive a single reward that is contingent upon a partial sequence or a complete trajectory. In this work, we define this challenging problem as Reinforcement Learning from Bagged Reward (RLBR), where sequences of data are treated as bags with non-Markovian bagged rewards. We provide a theoretical study to establish the connection between RLBR and standard RL in Markov Decision Processes (MDPs). To effectively explore the reward distributions within these bags and enhance policy training, we propose a Transformer-based reward model, the Reward Bag Transformer, which employs a bidirectional attention mechanism to interpret contextual nuances and temporal dependencies within each bag. Our empirical evaluations reveal that the challenge intensifies as the bag length increases, leading to the performance degradation due to reduced informational granularity. Nevertheless, our approach consistently outperforms existing methods, demonstrating the least decline in efficacy across varying bag lengths and excelling in approximating the original MDP’s reward distribution. The code is available at an anonymous link: https://anonymous.4open.science/r/RLBR-F66E/.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various domains, including autonomous driving Kiran et al. (2021), healthcare Yu et al. (2021), complex game playing Silver et al. (2016); Wurman et al. (2022), and financial trading Yang et al. (2020). One common and essential assumption for most RL algorithms is the availability of immediate reward feedback at each time step of the decision-making process. However, this assumption is violated in many real-world applications. Recognizing this gap, numerous studies Watkins (1989); Ke et al. (2018); Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023b) have explored the concept of delayed rewards, primarily focusing on trajectory feedback where rewards are allocated at the end of a sequence. Similarly, sparse reward settings, where agents receive infrequent and instance-specific feedback, present significant challenges in the well-known exploration-exploitation trade-off Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022).
On the other hand, real-world applications, such as autonomous driving (see Fig. 1), often feature complex non-immediate reward structures that neither sparse rewards nor trajectory feedback can fully capture. Providing reward for every action is impractical, and focusing only on end goals ignores crucial aspects of the journey. Typically, rewards are linked to completing specific tasks or sequences of actions, rather than individual actions or the final objective Early et al. (2022); Gaon and Brafman (2020). However, previous studies (see Fig. 2) mainly focused on learning desirable policies with immediate rewards Sutton and Barto (1998); Vinyals et al. (2019), trajectory feedback Watkins (1989); Arjona-Medina et al. (2019), or sparse rewards where feedback is infrequent and tied only to the current instance Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022), tending to fail under such scenarios.
To address these challenges, we introduce Reinforcement Learning from Bagged Rewards (RLBR), which better aligns with real-world scenarios by considering the cumulative effect of a series of state-action pairs. In RLBR, we define sequences of state-action pairs as bags, each associated with a cumulative bagged reward. This framework includes both the traditional RL setting, where each bag only contains a single instance (the first line of Fig. 2), and the trajectory feedback setting, where a bag spans an entire trajectory (the third line of Fig. 2). Furthermore, RLBR offers the potential to reduce the labeling workload by lessening the frequency of reward annotations. However, this benefit is balanced by increased learning complexity due to the reduced granularity of information.
Figure 1: An example of the reward bag structure on an autonomous driving trajectory. Each segment of the driving sequence is evaluated and assigned a score by an evaluator, and the score for each sequence is integrated based on the performance at each step.
@@
Our research contributes to the field in several ways. First, we establish the RLBR framework as a general problem setting and connect it theoretically to traditional Markov Decision Processes (MDPs), elaborated in Section 3. In Section 4, we introduce a Transformer-based reward model designed to assign rewards to individual instances while capturing environmental dynamics. Additionally, we propose an algorithm that alternates between optimizing this reward model and the policy, thereby enhancing the effectiveness of both components. Finally, in Section 5, our experiments show that the performance of baseline methods drops as the bag length increases, indicating that the larger bag length will increase the learning difficulty. Furthermore, we experimentally demonstrate the superiority of our method through comparative performance analyses and validate the ability of the proposed model to mimic the reward distribution of the ground truth MDP, highlighting its contextual understanding and adaptability to environmental dynamics.
@@ 2 Related Work
RL with Trajectory Feedback. RL with Trajectory Feedback (RLTF), termed episodic rewards or delayed rewards in some works (see the third line in Fig. 2), has become increasingly prominent in many applications (Watkins, 1989; Arjona-Medina et al., 2019; Zhang et al., 2023b). A key approach to this challenge is reward redistribution, aiming to assign rewards to individual instances more effectively. Return Decomposition for Delayed Rewards (RUDDER) Arjona-Medina et al. (2019) used a return-equivalent formulation for precise credit assignment, with advancements incorporating expert demonstrations and language models Patil et al. (2022); Liu et al. (2019); Widrich et al. (2021). Iterative Relative Credit Refinement (IRCR) Gangwani et al. (2020) presented a uniform reward redistribution model based on equal contributions from all state-action pairs, while Randomized Return Decomposition (RRD) Ren et al. (2021) proposed a novel upper bound for return-equivalent assumptions, integrating return decomposition with uniform reward redistribution. Additionally, Han et al. (2022) modified RL algorithms to use sequence-level information, helping agents learn from broader structures and long-term outcomes. However, our focus is on the original, unobservable rewards within reward bags, not the aggregated bagged rewards. Our experiments (see Section 5) show that the method by Han et al. (2022) is ineffective for long sequences.
Following the previous methodology, we adopt a reward redistribution learning strategy to enhance policy learning in the context of reward bags. However, conventional reward redistribution methods fail to effectively extract information from bag-level rewards, as they did not take the structure of the bag into consideration. Our experimental results demonstrate that directly applying previous methods to the reward bag setting does not yield results as promising as those achieved by our algorithm.
Figure 2: Comparison of four different types of reward settings in RL. In traditional RL, a reward is given based on each instance Sutton and Barto (1998); Vinyals et al. (2019), whereas in sparse reward settings, only some instances receive rewards Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022). Both types of rewards are Markovian. Trajectory feedback and the proposed bagged reward both address non-Markovian situations; however, trajectory feedback provides only one reward signal for the entire trajectory Watkins (1989); Arjona-Medina et al. (2019), while bagged rewards can include multiple reward signals within the sequence.
RL with Sparse Rewards. The sparse reward setting (see the second line in Fig. 2) presents significant challenges due to infrequent feedback, making it difficult for agents to effectively explore the environment and discover successful strategies. To address this challenge, various methods have been developed to enhance exploration. Reward shaping strategies Ng et al. (1999); Hu et al. (2020); Tambwekar et al. (2019) added rewards to actions in a way that guides the agent towards better policies without altering the original reward function. Curiosity-driven methods Pathak et al. (2017); Sekar et al. (2020) encouraged agents to explore the environment by visiting unseen states, potentially solving tasks with sparse rewards. Additionally, curriculum learning in RL Florensa et al. (2018); Riedmiller et al. (2018) involved presenting an agent with a sequence of tasks with gradually increasing complexity, allowing the agent to eventually solve the initially given sparse reward task.
In contrast to sparse rewards, which are typically infrequent Markovian signals associated with individual state-action pairs, bagged rewards depend on the cumulative effect of sequences of state-action pairs and are non-Markovian, requiring context from the entire sequence to accurately assess the contribution of each instance.
Transformers for RL. Transformers, introduced by Vaswani et al. (2017), have proven effective in RL environments requiring high sample efficiency and generalization, such as StarCraft Vinyals et al. (2019); Zambaldi et al. (2019) and DMLab30 Parisotto et al. (2020). They have also been used for sequential modeling in offline RL Chen et al. (2021); Janner et al. (2021b) and as reward models in policy learning from offline datasets Kim et al. (2023). Luo et al. (2021) combined deep convolution transfer learning models and inverse RL for reward function acquisition, while Zhang et al. (2023a) transformed non-Markovian reward processes into Markovian ones, enhancing online interaction efficiency. For reward bags, where rewards are linked to sequences of state-action pairs, Transformers are ideal due to their ability to handle long-range dependencies. Our approach leverages a Transformer-based model to redistribute bag-level rewards into instance-level rewards, capitalizing on the Transformer’s proficiency in sequential data analysis and attention mechanisms.
@@ 3 Reinforcement Learning from Bagged Reward
In this section, we first provide preliminaries on RL with immediate rewards and trajectory feedback, and then formulate the RLBR problem with an extension of the traditional MDP, the Bagged Reward MDP (BRMDP). We further conduct a theoretical analysis of the relationship between the BRMDP and the traditional MDP, showing that solving RLBR problems not only relaxes the demands on reward acquisition but also guarantees finding the optimal policy even with limited reward information.
@@ 3.1 Preliminaries
We consider finite-horizon RL in this paper, which is traditionally modeled using a finite MDP, where rewards are promptly provided for each state-action pair (Sutton and Barto, 1998). This paradigm is encapsulated in a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathit{P},\mathit{r},\mu)$, with $s\in\mathcal{S}$ and $a\in\mathcal{A}$ as the sets of states and actions, $\mathit{P}$ as the state transition probability function, $\mathit{r}$ as the immediate reward function, and $\mu$ as the initial state distribution. The primary objective in this framework is to discover a policy $\pi=p(a|s)$ that maximizes the cumulative sum of rewards over a horizon length $T$:
@@
Distinct from this traditional approach, RLTF offers feedback only after a complete sequence of actions, or a trajectory (Arjona-Medina et al., 2019; Zhang et al., 2023b). A trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T-1},a_{T-1})\}$ includes $T$ state-action pairs, with a reward $R_{\mathrm{traj}}(\tau)$ that is the sum of latent immediate rewards $\sum_{t=0}^{T-1}r(s_{t},a_{t})$, but only the cumulative reward is observable at the end of the trajectory. Denoting by $\mathcal{T}(\pi)$ as the distribution of trajectories induced by $\pi,P,\mu$, the learning objective in RLTF is to maximize the expected trajectory-based reward:
$J_{\mathrm{traj}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\big{[}R_{\mathrm{ep}}(\tau)\big{]}.$ (2)
@@ 3.2 Problem Formulation
We formally define the setting of RLBR, which has a granularity of rewards in between the above two settings. First, we define the notion of bags, which are sub-pieces of complete trajectories. A trajectory $\tau$ is divided into several neighboring bags, and a bag of size $n_{i}$, which starts from time $i$, is defined as $B_{i,n_{i}}=\{(s_{i},a_{i}),\ldots,(s_{i+n_{i}-1},a_{i+n_{i}-1})\},0\leq i\leq i+n_{i}-1\leq T-1$. Afterward, we define the BRMDP to navigate the complexities of the aggregated non-Markovian reward signals:
@@ A BRMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathit{P},\mathit{R},\mu)$, where
  • $\mathit{R}$ denotes the reward function over bagged reward: $\mathit{R}(B_{i,n_{i}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})$.
@@
The essential properties of BRMDP are two-fold: 1) only bagged rewards can be obtained by the learner, not immediate rewards; 2) the bagged reward is the accumulated sum of immediate rewards within the bag. In the RLBR framework, a bag $B_{i,n_{i}}$ metaphorically aggregates individual rewards from a contiguous sequence of state-action pairs into a unified reward unit. A trajectory $\tau$ is a composite of a set of bags, denoted as $\mathcal{B}_{\tau}$, which ensures that each trajectory includes at least one reward bag. We further assume a bag partition function defined by the environment: $\mathcal{G}:\tau\rightarrow\mathcal{B}_{\tau}$, which is a task-dependent function for generating bags given an input trajectory. Consequently, the learning objective of the policy is to maximize the accumulated bagged rewards:
@@
3.3 Equivalence of Optimal Policies in the BRMDP and the Original MDP
In the BRMDP, agents are unable to directly observe the immediate reward linked to each state-action pair. Nonetheless, we aim for agents to optimize policies in environments where bagged rewards are accessible and to perform well on the original task. To unify the learning objectives of the BRMDP with those of the traditional MDP, we formulate a theorem that crystallizes this relationship. The significance of this theorem lies in its assertion that the optimal policy derived in a BRMDP setting is congruent with that of an MDP when the bagged rewards are considered as cumulative sums. This finding not only validates the theoretical underpinnings of our approach but also affirms its practical relevance across various RL paradigms.
@@ 3.3 Equivalence of Optimal Policies in the BRMDP and the Original MDP
Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag. In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
The detailed proof of Theorem 1 is deferred to Appendix B. By this theorem, we can see that optimizing a policy on a BRMDP is equivalent to optimizing on the original MDP. The crux of the matter is how to accurately redistribute the bagged rewards so that the policies learned under these distributed rewards are more close to the policies on the original MDP.
Figure 3: The illustration of the Reward Bag Transformer (RBT) architecture. The Causal Transformer is used for reward representation by processing sequences of input data consisting of state-action pairs. The bidirectional attention layer is used for reward redistribution, utilizing the outputs of the Causal Transformer to predict instance-level rewards. Predicting the next state helps the model understand the environment, thereby improving reward prediction.
4 Reward Bag Transformer
This section delves into the intricacies of leveraging the bidirectional attention mechanism Seo et al. (2016); Vaswani et al. (2017); Devlin et al. (2019) for reward redistribution, designs a Transformer architecture, and outlines a comprehensive algorithm for the cultivation of efficient policies under the BRMDP framework.

### v3

@@ Abstract
In Reinforcement Learning (RL), it is commonly assumed that an immediate reward signal is generated for each action taken by the agent, helping the agent maximize cumulative rewards to obtain the optimal policy. However, in many real-world scenarios, designing immediate reward signals is difficult; instead, agents receive a single reward that is contingent upon a partial sequence or a complete trajectory. In this work, we define this challenging problem as RL from Bagged Reward (RLBR), where sequences of data are treated as bags with non-Markovian bagged rewards, leading to the formulation of Bagged Reward Markov Decision Processes (BRMDPs). Theoretically, we demonstrate that RLBR can be addressed by solving a standard MDP with properly redistributed bagged rewards allocated to each instance within a bag. Empirically, we find that reward redistribution becomes more challenging as the bag length increases, due to reduced informational granularity. Existing reward redistribution methods are insufficient to address these challenges. Therefore, we propose a novel reward redistribution method equipped with a bidirectional attention mechanism, enabling the accurate interpretation of contextual nuances and temporal dependencies within each bag. We experimentally demonstrate that our proposed method consistently outperforms existing approaches. The code is available at an anonymous link: https://anonymous.4open.science/r/RLBR-F66E/.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various domains, including autonomous driving Kiran et al. (2021), healthcare Yu et al. (2021), complex game playing Silver et al. (2016); Wurman et al. (2022), and financial trading Yang et al. (2020). One common and essential assumption for most RL algorithms is the availability of immediate reward feedback at each time step of the decision-making process. However, designing such immediate reward feedback is quite difficult and may mislead the policy learning in many real-world applications Kwon et al. (2023); Lee et al. (2023). Recognizing this gap, numerous studies Watkins (1989); Ke et al. (2018); Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023) have explored the concept of RL with Trajectory Feedback (RLTF), primarily focusing on trajectory feedback where rewards are allocated at the end of a sequence.
However, real-world applications, such as autonomous driving (see Fig. 1), often feature complex non-immediate reward structures that RLTF cannot fully capture. Providing rewards for every action is impractical, while focusing only on end goals ignores crucial aspects of the journey. Typically, rewards are linked to completing specific tasks or sequences of actions, rather than individual actions or the final objective Early et al. (2022); Gaon and Brafman (2020). Previous studies focused on learning desirable policies with immediate rewards (Sutton and Barto, 1998; Vinyals et al., 2019) or trajectory feedback (Arjona-Medina et al., 2019) are not ideal methods for such scenarios.
To address these challenges, we introduce RL from Bagged Rewards (RLBR), a framework that better aligns with real-world scenarios by considering the cumulative effect of a series of instances. In RLBR, sequences of instances are defined as bags, each associated with a bagged reward. This framework encompasses both the traditional RL setting, where each bag contains only a single instance, and the trajectory feedback setting, where a bag spans the entire trajectory, as special cases. Furthermore, RLBR offers the potential to reduce the labeling workload by lessening the frequency of reward annotations. However, this benefit is balanced by increased learning complexity due to the reduced granularity of information.
Figure 1: An example of the reward bag structure on an autonomous driving trajectory. Each segment of the driving sequence is evaluated and assigned a score by an evaluator.
@@
Our research contributes to the field in several ways. First, we establish the RLBR framework as a general problem setting by formulating it with an extension of the traditional Markov Decision Processes (MDPs), the Bagged Reward MDPs (BRMDPs), as elaborated in Section 3. In Section 4, we theoretically connect MDPs after reward redistribution to the original BRMDPs, which motivates us to employ a bidirectional attention mechanism within a Transformer-based framework to assign rewards to individual instances while capturing environmental dynamics. Additionally, we propose an algorithm that alternates between optimizing this reward redistributing model and the policy, thereby enhancing the effectiveness of both components. Finally, in Section 5, our experiments show that the performance of baseline methods drops as the bag length increases, indicating that the larger bag length will increase the learning difficulty. Furthermore, we experimentally demonstrate the superiority of the proposed method through comparative performance analyses and validate the ability of the proposed model to mimic the reward distribution of the ground truth MDP, highlighting its contextual understanding and adaptability to environmental dynamics.
@@ 2 Related Work
With the growing attention on RLTF, where trajectory feedback is often considered the sum of rewards over the entire trajectory, new methods specifically designed for RLTF have emerged. Return Decomposition for Delayed Rewards (RUDDER) Arjona-Medina et al. (2019) used a return-equivalent formulation for precise credit assignment, with advancements incorporating expert demonstrations and language models Liu et al. (2019); Widrich et al. (2021); Patil et al. (2022). Both Iterative Relative Credit Refinement (IRCR) Gangwani et al. (2020) and Randomized Return Decomposition (RRD) Ren et al. (2021) assumed that each state-action pair contributes equally to the trajectory feedback. Specifically, IRCR presented a uniform reward redistribution model, while RRD proposed a novel upper bound for return-equivalent assumptions, integrating return decomposition with uniform reward redistribution. Additionally, Han et al. (2022) modified RL algorithms to use sequence-level information, helping agents learn from broader structures and long-term outcomes. On the other hand, some studies on sparse rewards, such as reward shaping Ng et al. (1999); Tambwekar et al. (2019); Hu et al. (2020) and intrinsic reward methods Sorg et al. (2010); Pathak et al. (2017); Zheng et al. (2018, 2021), could be directly applied to RLTF. However, these approaches were primarily focused on addressing the exploration-exploitation trade-off in sparse reward settings, which made them less ideal for solving RLTF problems. A detailed discussion on the sparse reward setting and methods is provided in “Discussion on Sparse Rewards” in Appendix.
However, previous RLTF works that assume each state-action pair contributes equally to the trajectory feedback did not consider that instances contributing more to the feedback should receive greater attention. On the other hand, methods that consider sequence information struggled to capture effective information from long trajectories, resulting in suboptimal training outcomes, as evidenced by prior studies Ren et al. (2021); Zhang et al. (2023). Building on previous methods, we adopt a reward redistribution learning strategy to enhance policy learning in the context of reward bags. The proposed method aims to extract information from bag-level rewards, taking into account the contribution of each instance, ultimately learning an effective policy. Experimental results (see Section 5) demonstrate that the proposed method can effectively capture information from long sequences, and the rich information from the bags makes it more effective.
@@ 3 Reinforcement Learning from Bagged Reward
In this section, we first provide preliminaries on RL with immediate rewards and trajectory feedback, and then formulate the RLBR problem with an extension of the traditional MDPs, the BRMDPs.
@@ 3.1 Preliminaries
In traditional RL problems, which are modeled using finite MDPs, each state-action pair is promptly associated with a reward (Sutton and Barto, 1998). This paradigm is encapsulated in a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathit{P},\mathit{r},\mu)$, with $s\in\mathcal{S}$ and $a\in\mathcal{A}$ as the sets of states and actions, $\mathit{P}$ as the state transition probability function, $\mathit{r}$ as the immediate reward function, and $\mu$ as the initial state distribution. The primary objective in this framework is to discover a policy $\pi=p(a|s)$ that maximizes the cumulative sum of rewards over a horizon length $T$:
@@
Given that associating each state-action pair with a reward is challenging in real-world scenarios and involves significant labeling costs, RLTF, also known as episodic or delayed rewards in some works, has become increasingly prominent in many applications (Watkins, 1989; Arjona-Medina et al., 2019; Zhang et al., 2023). Distinct from traditional RL, RLTF offers only one feedback after a complete trajectory (Arjona-Medina et al., 2019; Zhang et al., 2023). A trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T-1},a_{T-1})\}$ consists of $T$ state-action pairs, with a cumulative reward $R_{\mathrm{traj}}(\tau)$ that is the sum of latent immediate rewards $\sum_{t=0}^{T-1}r(s_{t},a_{t})$, observable only at the end. Denoting by $\mathcal{T}(\pi)$ as the distribution of trajectories induced by $\pi,P,\mu$, the learning objective in RLTF is to maximize the expected trajectory-based reward:
$J_{\mathrm{traj}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\big{[}R_{\mathrm{traj}}(\tau)\big{]}.$ (2)
@@ 3.2 Problem Formulation
We formulate the RLBR problem as an extension of the traditional MDP, the BRMDP, which has a granularity of rewards in between the above two settings. First, we define the notion of bags, which are sub-pieces of complete trajectories. A trajectory $\tau$ is divided into several neighboring bags, and a bag of size $n_{i}$, which starts from time $i$, is defined as $B_{i,n_{i}}=\{(s_{i},a_{i}),\ldots,(s_{i+n_{i}-1},a_{i+n_{i}-1})\},0\leq i\leq i+n_{i}-1\leq T-1$. Afterward, we define the BRMDP to navigate the complexities of the aggregated non-Markovian reward signals:
@@ A BRMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathit{P},\mathit{R},\mu)$, where
  • $\mathit{R}$ denotes the bagged reward function, which define the reward over a bag: $\mathit{R}(B_{i,n_{i}})$.
@@
In the RLBR framework, a bag $B_{i,n_{i}}$ metaphorically a unified reward unit for a contiguous sequence of state-action pairs. A trajectory $\tau$ is a composite of a set of bags, denoted as $\mathcal{B}_{\tau}$, which ensures that each trajectory includes at least one reward bag. We further assume a bag partition function defined by the environment: $\mathcal{G}:\tau\rightarrow\mathcal{B}_{\tau}$, which is a task-dependent function for generating bags given an input trajectory. Consequently, the learning objective of the policy is to maximize the accumulated bagged rewards:
@@
Figure 2: The illustration of the Reward Bag Transformer (RBT) architecture. The Causal Transformer is used for reward representation by processing sequences of input data consisting of state-action pairs. The bidirectional attention layer is used for reward redistribution, utilizing the outputs of the Causal Transformer to predict instance-level rewards. The next state prediction helps the model understand the environment, thereby improving reward prediction.
4 Reward Redistribution for RLBR
This section delves into the proposed reward redistribution method for RLBR. We aim for the reward redistribution to meet the following criteria: first, the policies learned from the MDP after reward redistribution should be consistent with those learned from the BRMDP; second, the redistribution process should reflect each instance’s contribution based on the sequence information. The subsequent subsections are structured to first theoretically prove the rationality of reward redistribution, followed by proposing the use of a bidirectional attention mechanism for this redistribution, and finally outline a comprehensive algorithm for developing efficient policies within the BRMDP framework.
4.1 Equivalence of Optimal Policies in the BRMDP and the MDP after Reward Redistribution
In the BRMDP, agents do not have direct access to the immediate rewards associated with each instance. To address this, we adopt a common approach used in trajectory feedback scenarios, modeling the cumulative sum form as the reward redistribution strategy to convert bagged rewards into instance-level rewards Ren et al. (2021); Arjona-Medina et al. (2019); Zhang et al. (2023). Specifically, we define $\mathit{R}(B_{i,n_{i}})=\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t})$, where $\hat{r}(s_{t},a_{t})$ represents the redistributed reward. However, a critical question arises: How can we ensure that the policy learned with redistributed rewards performs well in the original BRMDP? To address this, we provide the following theorem:
@@ 4.1 Equivalence of Optimal Policies in the BRMDP and the MDP after Reward Redistribution
Provide a BRMDP where the bagged reward is regarded as the sum of rewards for state-action pairs of a MDP after reward redistribution contained within the same bag. In this context, the set of optimal policies for the BRMDP $\Pi_{\mathrm{B}}$ aligns with that of the MDP after reward redistribution $\Pi$, implying that $\Pi=\Pi_{\mathrm{B}}$.
The detailed proof of Theorem 1 is deferred to the “Omitted Proof” in Appendix. By this theorem, we can see that optimizing a policy on a MDP after reward redistribution is equivalent to optimizing on the original BRMDP.